Write the slope-intercept equation for the line that passes through (10, -4) and has slope -3/10. Please show all of your work.
Answer by rfer(16322)

y=mx+b
-4=-3/10*10+b
-4=-3+b
-1=b
--------------
y=-3/10x-1